Ganymede is one of the moons of Jupiter.Ganymede's period of revolution around Jupiter is 7.16 earth days and orbital radius about 1.7x10^9m.Find the mass of Jupiter.


1
Expert's answer
2021-12-01T10:28:52-0500

Explanations & Calculations


  • Apply Newton's gravitation law to Saturn and its moon.
  • An earth day is "\small 86400" seconds.
  • Then,

"\qquad\qquad
\begin{aligned}
\small G\frac{M_S.m_G}{r^2}&=\small m_Gr\omega^2\\
\small M_S&=\small \frac{r^3}{G}.\frac{4\pi^2}{T^2}\\
&=\small \frac{(1.7\times10^9\,m)^3}{6.67\times10^{-11}m^3kg^{-1}s^{-2}}.\frac{4\pi^2}{(7.16\times86400\,s)^2}\\
&=\small 7.60\times10^{27}
\,kg\end{aligned}"